OpenStudy (anonymous):

If x is a whole number and 437 = (21 + x)(21 - x), then x =

OpenStudy (wolf1728):

437 = (21 + x)(21 - x) -x^2 +441 = 437 -x^2 +878 = 0 a=-1 b=0 c=878 Putting this into quadratic formula: x= 0 +- square root (0 -4*-1*878)/2*-1 x= +- square root ( -4*-1*878)/-2 x = +-square root (3,512) / -2 x = +-square root (878) x = +29.6310647801 x = -29.6310647801

Directrix (directrix):

441 - 4 = 437 (Difference of 2 squares) 441 - 4 = (21 + 2) * (21 - 2) = 437 What is the value of x? @kbriones0

OpenStudy (anonymous):

437?

OpenStudy (wolf1728):

kbriones0, the value of x has already been posted by me I used the quadratic formula (scroll back just a bit)

Directrix (directrix):

I got the value of x as being 2. @wolf1728 I am not following how the answer of x = +29.6310647801 x = -29.6310647801 was derived.

Directrix (directrix):

@wolf1728 >> x is a whole number

OpenStudy (wolf1728):

I made a mistake in my previous answer: 437 = (21 + x)(21 - x) -x^2 +441 = 437 That becomes -x^2 = -4 x^2 -4 = 0 (x+2)*(x-2) =0 x = 2 x = -2